5x(4th power)(2xy-9x(cubed))
Answer by nyc_function(2733)

(5x^4) times (2xy - 9x^3) =

Multiply (5x^4) by every term inside and add the exponents.

(5x^4) times (2xy) = 10x^5(y).

(5x^4) times (-9x^3) = -45x^7

We now put it all together.

10x^5(y) - 45x^7

Done!